VIBRATION TO ELECTRONIC ENERGY TRANSFER IN VIBRATION-VIBRATION PUMPED CARBON MONOXIDE ^{a}

\underline{E. PLÖNJES ^{b}}, I. ADAMOVICH, J.W. RICH, Department of Mechanical Engineering, The Ohio State University, Columbus, OH 43210.

In a CO-Ar gas mixture carbon monoxide is excited into low vibrational states v=1,...,10 by absorption of CO laser radiation. Vibration-vibration exchange pumping results in the diffusion of energy into very high vibrational states. Yet states above v = 40 have never been observed to be populated. As v = 40 is isoenergetic with the A^{1}Π state we believe this effect to be due to vibration to electronic energy transfer. A master equation kinetic model which takes into account vibration-vibration, vibration-translation, vibration-electronic energy transfer processes and radiative decay has been developed and will be compared to measurements of kinetic rates for the vibrational to electronic transfer. The diffusion of energy through the vibrational quantum states is observed by time resolved step-scan FTIR spectroscopy on the Δv = 2 emission with spectral resolution of 8 cm^{−1} and time resolution of 5μs. Vibrational to electronic energy transfer is monitored using step-scan UV Fourier transform spectroscopy on the 4th positive emission of CO.

aSupport by AFOSR Space Propulsion and Power Program under Grant No. F49620-96-1-0184 for the preparation of this work is gratefully acknowledged

bVisiting Scholar at The Ohio State University, permanently at the Institut für Angewandte Physik der Universität Bonn, 53115 Bonn, Germany